Solve for P: 7(3P+4) = 8(2P+5) + 13
Please help me solve for P...?
Am I correct with 21P=4=16P+5+13?
Thank You


Answer by scianci(186)

7(3P+4) = 8(2P+5) + 13
Distribute:
21P + 28 = 16P + 40 + 13
solve for P in the usual manner